OpenStudy (anonymous):

Bayside insurance offers two health plans. Under plan A,Giselle would have to pay the first $190 dollars of her medical bills, plus 30% of the rest. Under plan B, Giselle would have to pay the first $210, but only 25% of the rest. For what amount medical bills will plan B save Giselle Money? Assume she has over $210 in bills. Giselle would save with plan B if she had more than $____ in bills. (round to the nearest whole number)

OpenStudy (amistre64):

190 +(bill-190)(.3) 210 +(bill - 210)(.25)

OpenStudy (amistre64):

A=bill(.3) +190(.7) B=bill(.25) + 210(.75)

OpenStudy (anonymous):

Plant A=0.3x+190 Plant B=0.25x+210 0.25x+210=0.3x+190 210-190 =0.3x-0.25x 10=0.05x 10/0.05=x 200 =x

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

She have to spent $200 amount medical bills will save on plan B

OpenStudy (amistre64):

A=bill(.3) +133 B=bill(.25) + 157.50 A-B = 0 bill(.3-.25) +133-157.50 = 0 bill(.05) -24.50 = 0 bill = 24.50/.05 = 490

OpenStudy (amistre64):

490(.30) + 133 = 280 490(.25) + 157.50 = 280 ... anything over 490 she saves with plan b

OpenStudy (anonymous):

Let x equal the total medical expenses. Equate the cash payouts for each plan and solve for x.\[190+\left(\frac{3}{10}\right)(x-190)\text{=}210+\left(\frac{1}{4}\right)(x-210) \]x = 490, the same as amistre64.